After 50 times at bat, a major league baseball player has a batting average of 0.160. How many consecutive hits must the player get to raise his batting average to 0.250?

i tried solving this problem this way:
0.250 - 0.160 = 0.09
50/0.09 = 555.555....

Answer by stanbon(75887)

After 50 times at bat, a major league baseball player has a batting average of 0.160. How many consecutive hits must the player get to raise his batting average to 0.250?
-----------------------
After 50 times at bat he has 0.16*50 = 8 hits
-----------------
Let x be the number of hits he needs to be batting 0.250
EQUATION:
# hits/# at bat = (x + 8)/(50+x) = 0.250
x + 8 = (1/4)(50+x)
4x + 32 = 50 + x
3x = 18
x = 6 (# of consecutive hits he needs to be batting 0.250)
============================
Cheers,
Stan H.